OpenStudy (anonymous):

A car is traveling around a horizontal circular track with radius r = 190.0 m as shown. It takes the car t = 58.0 s to go around the track once. The angle θA = 34.0° above the x axis, and the angle θB = 61.0° below the x axis. What is the x/y acceleration component at position b?

OpenStudy (anonymous):

Vertical Component: \[a_y = a(\sin61)\] Horizontal Component: \[a_x = a(\cos61)\] \[(a)Centripetal Acceleration= v^2/r\]\[(v)Velocity = Distance/Time\] \[\pi *2(190)=1193.80m , 1193.80/58= Velocity :: 20.58m/s^2 \] \[(a)Centripetal Acceleration = (20.58m/s^2)^2/190m= 2.23 m/s^2\]